# Finite solvable groups of square-free Euler characteristic

### v1

@@ 4. $p$-Groups
Let $(x,y,z)$ be a regular triple for $G$, and let $\overline{G}=G/N$ for a proper normal subgroup $N\lhd G$. If $\overline{x},\overline{y},\overline{z}$ are all involutions of $\overline{G}$ with $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$; otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$, while in both cases, $\overline{G}$ is generated by one or two involutions, namely, $G$ is ${\sf Z}_{2}$ or dihedral, respectively. $\Box$
@@
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$, and $K/H\cong{\rm S}_{3}$. Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$. Recall that $(x,y,z)$ is a regular triple for $G$. Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
@@
Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$. By Lemma 2.1, each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index, and so G=(A:B):KG=(A{:}B){:}K as described in Theorem 1.1.
Note that $G$ has a regular triple. Then it is of even order, and so $K\neq 1$. Further, if $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$, then $K$ has $G_{2}={\sf Z}_{2}^{3}$ as a minimal normal subgroup. By checking the $p$-groups appearing in Lemma 4.3, meanwhile noticing that ${\sf Z}_{2}^{3}$ cannot be embedded into ${\rm GL}(2,p)$ for all primes $p$ (as a faithful representation of ${\sf Z}_{2}^{3}$ on $\mathbb{F}_{p}$ is of degree at least $3$), we deduce that $G_{2}={\sf Z}_{2}^{3}$ centralizes each Sylow subgroup of A:BA{:}B. Hence $G_{2}\lhd G$, and so $G$ has an odd factor group $G/G_{2}$, leading to a contradiction by Lemma 6.1. Therefore, $K=G_{2}$ or $G_{\{2,3\}}$.
Suppose $K=G_{2}$. Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple and then it is one of the groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$ or a dihedral group. In conclusion, $K$ is ${\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1$.
Suppose $K=G_{\{2,3\}}$, which is one of the groups given in Lemma 5.6 (ii) - (iii). Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple, or $K$ is a dihedral group. Further, note that a factor group of a dihedral group is also dihedral (or ${\sf Z}_{2}$). Thus, $K$ has no odd factor groups, and has no factor groups isomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ either, because by Lemma 6.2 neither ${\rm Q}_{8}{:}{\rm S}_{3}$ nor ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple, and meanwhile, neither of them is a dihedral group. Then, by checking the groups appearing in Lemma 5.6 (ii) - (iii), we find that the only candidates for $K=G_{\{2,3\}}$ are the following groups in Table 1: $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$ and the two groups in Table 2: $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$. $\Box$

### v2

@@ 4. $p$-Groups
Let $(x,y,z)$ be a regular triple for $G$. If $\overline{x},\overline{y},\overline{z}\in\overline{G}$ are all involutions, and $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$. Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$. In both cases, $\overline{G}$ is generated by one or two involutions and then it is ${\sf Z}_{2}$ or dihedral.
It is similar to prove the remaining part of this lemma. $\Box$
@@
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$ and $K/H\cong{\rm S}_{3}$. Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$. Recall that $(x,y,z)$ is a regular triple for $G$. Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
@@
Lemma 6.3.
Let $G=({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, with $a^{c}=a^{-1}$, $b^{c}=b^{-1}$. Then $G$ has no rotary pair.
Proof.
Let $(\alpha,z)$ be a rotary pair for $G$. Since $G$ is not dihedral, $|\alpha|\neq 2$, so $\alpha\in\langle a,b\rangle$ and $\langle\alpha\rangle$ is normal in $G$. It follows that $\overline{G}=G/\langle\alpha\rangle=\langle\overline{z}\rangle={\sf Z}_{2}$, not possible. $\Box$
@@
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be a solvable arc-transitive automorphism group of $\mathcal{M}$}. By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so that
G=(A:B):KG=(A{:}B){:}K
as described in Theorem 1.1.
Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties: (i) $G$-regular; (ii) $G$-vertex-reversing; (iii) $G$-vertex-rotary; respectively in each case, the group $G$ has: (i) a regular triple; (ii) a reversing triple; (iii) a rotary pair. In particular, $G$ is of even order, and so $K\neq 1$.
First, suppose $K=G_{2}$. By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ with $\ell\geqslant 2$, while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
Then, suppose $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$. In this case, $G$ has an odd factor group $G_{7}$ or $G_{7}{:}G_{3}$, respectively. By Lemma 6.1, G has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic. It follows that $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$.
At last, suppose that $K=G_{\{2,3\}}$ is one of the groups given in 5.6 (2) - (3).
(a) If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1. Hence, $K$ is one of the groups in Tables 1 - 2, excluding those corresponding to cases (1.3), (1.4), (1.7), (2.1) - (2.3) as they have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$.
(b) If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple. In particular, $G$ has a reversing triple, and so $K$ is one of the groups in (a). Furthermore, by Lemma 6.2 the groups which have ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ as a factor group should be excluded. It follows that $K$ is one of: $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where
$X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
(c) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, with each odd factor group cyclic by Lemma 6.1. So, either $K=X{:}K_{3}=X{:}{\sf Z}_{3^{\ell}}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$ or $K$ is one of the groups in Tables 1 - 2. For the former, $X\neq{\rm Q}_{8}$, as the only involution of ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ lies in the center and so it does not have a rotary pair. For the latter, by Lemma 6.3 the groups corresponding to cases (1.5) and (1.6), which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded.
In conclusion, $K$ is one of the groups stated in Theorem 1.2. $\Box$